### v1

@@ 6.3. Translucent Selective Deanonymizations
We now discuss how we use a translucent ledger to create a translucent deanonymization process. After identifying suspicious Seldom-anonymized traffic, a LEA requests a deanonymization from the consortium. Therefore, the consortium first assesses this request through a translucent voting. If the request is accepted, the consortium, with the help of the exit relay, peels away the layers of threshold encryption from the client’s encrypted identity, such that only the requesting LEA learns the identity of the suspect.
@@ 6.3.1. Translucent Voting
We assume that LEA monitor the Internet or request logs from websites to identity suspicious Seldom-anonymized traffic for which knowledge about the origin would benefit public safety. The LEA then first searches the state-sponsored database for the flow record in question. For now, consider that this database is fully accessible to LEA, such that LEA have a list of all communication flows exiting Seldom. To further strengthen privacy, this database could be based on searchable encryption (and rate-limited through cryptographic puzzles or the database being operated by the consortium) where each exit relay logs to a different index with an encryption keys only known to the relay. This index could still be searched by LEA to match concrete flow records, while no information about all traffic leaks.
After identifying the appropriate flow record, the LEA can request a deanonymization from the consortium, which subsequently assesses the request through a voting process, also logging the request and result to the translucent ledger. Appendix D shows a examplary smart contract used for this task for a consortium with five members and basic majority voting. The initializing of the request as well as the finalizing of a vote will trigger events in the managing smart contract and are thus revealed to the public immediately. The delayed disclosure ensures that the concrete request, the reasoning behind it, and the vote’s result is eventually fully disclosed. Following a successful vote, only the exit relay can decrypt the key in the flow record linking it to the encrypted identity.
@@ 6.3.2. Linking Flows to Encrypted Identities
To link the identified flow record to an encrypted identity, the consortium sends the record to the exit relay from which the traffic originated. With the request, the consortium also provides a Merkle proof of the accepted request. For this proof, the consortium reveals the smart contract’s variables storing the digest of the voted-upon request and the variable storing the vote’s outcome. Only the exit relay is informed about this request, as Seldom would otherwise extensively interfere with ongoing investigations. After verifying these proofs, the exit relay is thus sure that the consortium agreed on the necessity of this deanonymization and that this is eventually disclosed. Then, the exit relay decrypts the database key as well as the flow record signature by the client identification key. The exit relay then transfers this data to the consortium.
The consortium can now retrieve the encrypted identity associated with the database key by jointly threshold-decrypting this encrypted identity, before verifying the honesty of the exit relay. First, the data, pointed to by the key, has to be retrievable. Second, the signature of the flow record must be verifiable with the client identification key revealed by the decryption. This signature ensures that the outer encrypted identity is indeed the origin of the investigated communication, as only the client knows the corresponding private key and would not sign flow records of unknown connections. Third, the consortium has to check that the indicated predecessor relay has a known identity.
Finally, the consortium ensures that this middle relay did indeed vouch for the correctness of the next layer’s encrypted identity by verifying . Knowing the middle relay’s identity means that either a valid certificate is provided, or, more commonly, the relay’s identity is stored in Seldom’s consensus. Together, these verification steps convince the consortium that the exit received the claimed inner encrypted identity from the indicated predecessor relay, and that the exit relay acted according to the protocol. If any of the preceding verifications fail, this would indicate a malicious exit relay and thus likely collusion with the client.
@@ 6.3.3. Extracting the Traffic’s Origin
After verifying the exit relay’s honesty, the consortium proceeds to similarly verify the other relay’s honesty, ultimately leading to the identification of the client’s identity. Therefore, the consortium first jointly decrypts the inner encrypted identity . Then, the consortium verifies the honesty of the middle relay with the same process as for the exit relay. Additionally, the consortium verifies that the hop key matches that of the outer encrypted identity layer and that this key is correctly signed by . Again, if any verification fails, the middle relay behaved maliciously, and the consortium reports this as likely collusion with the client to the LEA.
If a circuit consists of more than three relay’s the exact same procedure can be used to peel additional layers from the encrypted identity. Only the last layer is treated differently. As the last threshold decryption reveals the potential origin of the traffic, the resulting decryption shares are only communicated to the requesting LEA. Thus, no member of the consortiums learns this identity. This restriction maximizes privacy as only the requesting LEA learns the client’s identity, while the consortium only knows the relays.
The LEA then still has to verify the correctness of the decrypted client identity to ensure that the entry relay did not manipulate this information. Therefore, the LEA first verifies that the IP certificate is valid and that a trusted CA issued it. Additionally, the LEA verifies the correctness of the hop key , analogously to this verification by the consortium as previously discussed. To this end, the consortium shares the hop key claimed in the previous layer with the LEA. If both checks succeed, the revealed identity is indeed that of the user from which the suspicious traffic originated. Otherwise, the entry relay attempted to cover up the true identity, and the LEA should investigate this. Through Seldom’s iterative process, the requested identity is thus only revealed to the requesting LEA, while misbehavior (i.e., sending no or false data) during the anonymization process can be clearly attributed to one relay.
@@ 6.4. Probing Relay Honesty
Incorrectly behaving relays are revealed during deanonymizations. However, such misbehavior could not only occur from collusion with the client, but also from genuine errors or general malice to damage the network. Therefore, the relay’s protocol abidance under normal circumstances should be verified without risking individual’s privacy, such that misbehaving nodes can be fixed or removed from the network. Regular probing by the directory authorities also reduces the likelihood that an exit relay’s key loss in a critical situation is coincidental. This feature significantly strengthens our assumption that failures during deanonymization are due to a relay colluding with a client, warranting further investigation.
@@
To generate probe traffic for verifying the honesty of relays, Seldom relies on cryptographic commitments. Therefore, a directory authority, i.e., the prober, first generates a key later used as the client’s identification key during the probe. To create a cryptographic commitment of the probe, the prober hashes this key, the probed exit relay, a traffic destination, and a timeframe for the probe. The prober then stores this commitment in the translucent ledger. Afterwards, the prober establishes a new circuit with the probed relay as the last onion router and sends traffic as specified in the commitment and using the pre-generated client identification key. Afterwards, the prober reveals the preimage of the commitment on the ledger, and it requests an deanonymization. The consortium accepts this vote, and the exit relay is asked to initiate the deanonymization process.
@@
At this point, the exit relay does not know that the deanonymization request is only for probe traffic. It learns that the vote for the request was successful, but does not see the reasoning behind it which is only revealed during the delayed disclosure of the request to the public. The exit relay thus is expected to cooperate like for ordinary deanonymization. Afterwards, the consortium verifies the signature of the database key with the public key indicated in the commitment. If this verification is successful, the exit relay passed the probe. Otherwise, the directory authority no longer accepts this exit relay. The other directory authorities can adopt this assessment but are encouraged to verify it through independent probes. Otherwise, a single directory authority would have the power to remove relays at will by claiming they failed a probe. The consortium can also further investigate the encrypted identity to identify non-exit relays that are misbehaving. Through regular probing, the consortium and the LEA can be much more confident that each relay behaves correctly during deanonymization.
@@
Probing is expected to make up the bulk of all requests. As the public is informed about each request immediately, all probers are encouraged to reveal that a probe took place immediately after conclusion. Therefore, they can call a special function in the smart contract that triggers an event to inform the public, such that the public always knows how many requests actually deanonymize a user.
@@ 7.1. Evaluation Setup
Most new Tor features are evaluated through simulations. However, simulation abstracts from Tor’s cryptographic processing (Jansen and Hopper, 2012), which is exactly the part impacted by Seldom (and not the behavior of network). Thus, we evaluate Seldom in a local deployment. Since our changes do not impact the network as a whole, but rather individual relays or circuits, this still gives accurate results on Seldom’s overhead.
We modify Tor version 0.4.1.5 to implement Seldom. As our target host for connection establishments we run an apache2 web server. We additionally host a local CA for IP certificates and a MongoDB database to store deanonymization information. The consortium members’ behavior is expressed in Python and the translucent ledger is a permissioned Ethereum blockchain (Wood et al., 2019) modified to realize the translucency properties. Our server is equipped with two Intel® Xeon® Silver 4116 Processors (24 CPU cores/48 threads @ ) and of RAM. We run nine relays: one directory authority, three exit relays, and five non-exit relays. Each process (e.g., databases, Tor clients, Tor relays) is pinned to one CPU core to prevent cross-process interference.
@@ 7.2. Circuit Establishment
We compare the latency to extend a circuit by one hop at a time in Figure 5. The x-axis shows the elapsed time, while the different horizontal bars represent circuits of various lengths in Seldom or Tor respectively. The bars represents the total time to establish a circuit. We notice a processing overhead of for establishing the first hop, which increases to for the sixth hop. Overall, we observe a latency increase of to , depending on circuit length. The receipt of an IP certificate causes an overhead of . The remaining overhead stems from signature processing, threshold encryption, and transferring the partially encrypted identity to the next hop. All these steps require more time the more previous hops exist (as more data has to be processed). The observed variance in timings results from occasional failures of the path selection algorithm due to the small network size. In a global overlay network, additional latency in the range of multiple 100 milliseconds and jitter would be added by the network and make this overhead insignificant. Overall, Seldom thus does not perceivably impact circuit establishment.
@@ 7.3. Connection Establishment
Analyzing the impact of Seldom on connection establishment is essential as additional latency in this step is directly affecting user experience. To assess this overhead, we request a web page from the locally hosted web server through cURL, which uses socks5 to communicate with the Tor client. We measure the time to first byte, i.e., the time between issuing the request and receiving the first byte through the established circuit. We repeat this measurement for different circuit lengths (approx. 700 measurements each) and compare a vanilla Tor deployment to Seldom. Our results are shown in Figure 6. We observe that Seldom takes on average longer than Tor to establish a connection and receive the first byte of the response, independent of the circuit length. This delay stems from the generation and verification of the flow record signature as well as writing this flow record to a local database before the connection is established. Again, these delay are overshadowed by the latency of the overlay network which is two orders of magnitudes higher.
@@ 7.4. Bandwidth and Storage Requirements
While Seldom barely impacts performance, the additional data that is transferred and stored risks to become the real bottleneck. To quantify this overhead, we first measure how much bandwidth exit relays consume to transfer all deanonymization data to a state-sponsored database. During our evaluation, the client establishes circuits with a length of three and a fixed exit relay. A fixed number of connections is established to random IP addresses and ports through each circuit before it is discarded. The exit relay waits for a predefined amount of closed circuits before it migrates the combined local storage for these circuits to the state-sponsored database. When migrating, mongodump is executed on the local mongoDB database and the resulting folder is serialized and compressed using zlib. The resulting archive is sent over a TLS channel to a state-sponsored database. Simultaneously, the hash of this archive is sent to the translucent ledger to ensure consistency, across multiple governmental databases.
@@
In Figure 7, we show how bandwidth usage depends on the number of connections established per circuit and how often the local buffer is flushed. Flow records are approximately an order of magnitude smaller than encrypted identities, which are only generated for each circuit and not each connection. Furthermore, the more data is migrated to the database in one go, the more the local buffer can be compressed. However, the benefit of sending more data at once levels off quickly, indicating that frequent flushing is desirable. While using an established circuit for more connections is beneficial, such a change directly influences the susceptibility to deanonymization attacks. Recent measurements by Mani et al. estimate the average number of established connections per circuit to be about 20 (Mani et al., 2018). An exit relay that thus migrates its database after 100 circuits have been established that were used to establish 20 connections each, would produce (99%-CI:[; ]) of data per connection.
@@
The measurements of Mani et al. also allow us to predict the storage requirements of a Tor-sized Seldom network. For this prediction, we project the amount of generated deanonymization data based on the activity in the Tor network over time. According to Mani et al., an average of 2.1 billion (95%-CI:[1.7 billion; 2.5 billion]) exit streams were created every 24 hours in April 2018 (Mani et al., 2018). Over the same period, the Tor Project reported an average of of consumed bandwidth by the Tor network (The Tor Project, 2024). For our estimations, we assume a constant relationship between the amount of consumed bandwidth and the amount of exit streams over the last six years. This is a worst-case assumption as the size of web pages grew by over , while the number of TCP connections per web page shrunk by about over the same period (http archive, 2024). Thus, in practice, parts of the additional bandwidth consumed by Tor is due to larger web pages which are downloaded over fewer connections.
@@
Figure 8 thus quantifies the bandwidth overhead of Seldom in a Tor-sized network according to those estimations. Here, the blue line shows the global bandwidth consumed by all exit relays to write data into the state-sponsored database over time with confidence intervals (blue area). In January 2020, all exit relays combined would consume (95%-CI:[; ]) to write deanonymization data to the state-sponsored database. To put these numbers into perspective, this would mean a merely (95%-CI:[; ]) increase in data transmitted by exit relays compared to Tor. A Seldom exit relay advertising of exit traffic per second would thus insert approx. of deanonymization data into the database. This estimate also means that this same relay has to, on average, verify 385 ed25519 signatures and compute 18 hybrid threshold encryptions per second, a processing overhead hardly impacting exit relays. Figure 8 also highlights the amount of data that would be retained in the state-sponsored database, assuming a retention period of two years. The database would thus have to store (95%-CI:[; ]) of deanonymization data, which can easily be provided by government organizations.
@@
Seldom can thus easily scale to Tor’s current size. If Seldom were to significantly outgrow Tor, we sketch several scalability improvements. First, websites that are unlikely to be abused for illegal activities (e.g., news websites) could be whitelisted, i.e., exit relays would not store flow records (and encrypted identities if only such websites are visited). Secondly, Seldom could rely on probabilistic deanonymization, where only a randomly selected subset of identities is deanonymizable. Therefore, the consortium can retroactively select a fraction of records that are persistently stored through provable randomness. Thus, bandwidth and storage overhead are unlikely to limit Seldom’s scalability.

### v2

@@ 6.3. Translucent Selective Deanonymizations
We now discuss how we use a translucent ledger to create a translucent deanonymization process. After identifying suspicious Seldom-anonymized traffic, an LEA requests a deanonymization from the consortium. Therefore, the consortium first assesses this request through a translucent voting. If the request is approved, the consortium members, with the help of the exit relay, jointly remove the layers of threshold encryption from the client’s encrypted identity. The final decryption shares are then only shared with the requesting LEA, such that the identity of the suspect is not even revealed to the consortium.
@@ 6.3.1. Translucent Voting
We assume that LEA monitor the Internet or request logs from websites to identify suspicious traffic that has been anonymized by Seldom when deanonymization is warranted to uphold public safety. As a first step, the LEA then searches the database for the flow record (cf. Figure 3) that matches the observed traffic. For now, consider that this database is fully accessible to LEA, such that LEA have a list of all communication flows exiting Seldom. To further strengthen privacy, this database could be based on searchable encryption (Curtmola et al., 2006) (and rate-limited through cryptographic puzzles or the database being operated by the consortium) where each exit relay logs to a different index with an encryption key only known to the relay. This index could still be searched by LEA to match concrete flow records, while no information about all traffic leaks.
After identifying the appropriate flow record, the LEA can request a deanonymization (and supply evidence) from the consortium by calling the corresponding function of the smart contract deployed on the translucent blockchain. This deanonymization request contains one suspicious flow record (referenced by its hash) and justification for the consortium to approve the request, which are both written to the translucent ledger. What would be accepted as valid justification must be agreed upon by the consortium, but it could involve a warrant or the evidence collected by the LEA. Appendix D shows an exemplary smart contract for a consortium with five members and basic majority voting. This function call triggers an event in the smart contract which reveals that a deanonymization request occurred immediately to the public, while the suspicious flow record and justification will only later be revealed to the public.
Then, each consortium member can call the smart contract to vote in favor of the deanonymization. If enough positive votes are collected, the request is accepted; this approval triggers an on-chain event visible to all Seldom users, i.e., the public. The delayed disclosure then ensures that the individual votes of each consortium member will eventually be fully published. This event also starts the deanonymization procedure by the consortium. Following a successful vote, however, only the exit relay can decrypt the key in the flow record linking it to the encrypted identity.
@@ 6.3.2. Linking Flows to Encrypted Identities
The consortium needs to request the exit relay to link the identified flow record to an encrypted identity. Therefore, the exit relay needs to be informed about the flow record and convinced that the request was approved and logged by the consortium. The consortium provides this proof through an EIP1186 Merkle proof (Jentzsch and Jentzsch, 2018), which proves the deanonymization request (without justification) and voting record to the exit relay by revealing the hash chain that connects the smart contract state variables to the already public blockchain header. This communication between the consortium and the exit relay is confidential and realized over a TLS channel between one consortium member and the exit relay.
Thus, only the exit relay is informed about this request, as Seldom would otherwise extensively interfere with ongoing investigations. After verifying the request, the exit relay is thus sure that the consortium agreed on the necessity of this deanonymization and that this will eventually be disclosed. Then, the exit relay decrypts the digest , which is the database key for the encrypted client identity , as well as the flow record signature by the client identification key. The exit relay then transfers this data to the consortium.
The consortium can now retrieve the associated encrypted identity (cf. Figure 2) and remove the first threshold-encryption layer to verify the honesty of the exit relay. For this verification, several requirements must be fulfilled. First, the data pointed to by the database key has to be retrievable. Second, the signature of the flow record must be verifiable with the client identification key revealed by the decryption. This signature ensures that the outer encrypted identity is indeed the origin of the investigated communication, as only the client knows the corresponding private key and would not sign flow records of unknown connections. Third, the consortium has to check that the indicated predecessor relay has a known identity . Knowing the middle relay’s identity means that either a valid certificate is provided or, more commonly, the relay’s identity is stored in Seldom’s consensus. Finally, the consortium ensures that this middle relay did indeed vouch for the correctness of the next layer’s encrypted identity by verifying the signature .
Together, these verification steps convince the consortium that the exit relay received the claimed inner encrypted identity from the indicated predecessor relay and that the exit relay acted according to the protocol. If any of the preceding verifications fail, this would indicate a malicious exit relay and thus likely collusion with the client.
@@ 6.3.3. Extracting the Traffic’s Origin
After verifying the exit relay’s honesty, the consortium proceeds to similarly verify the other relays’ honesty, ultimately leading to the identification of the client’s identity. Therefore, the consortium first jointly decrypts the inner encrypted identity . Then, the consortium verifies the honesty of the middle relay with the same process as for the exit relay. Additionally, the consortium verifies that the hop key matches that of the outer encrypted identity layer and that this key is correctly signed by . Again, if any verification fails, the middle relay behaved maliciously, and the consortium reports this as likely collusion with the client to the LEA. Otherwise, the middle relay acted according to the Seldom protocol.
If a circuit consists of more than three relays, the same procedure can be used to iteratively unveil further intermediate layers of the encrypted identity. Only the last layer is treated differently. While the encrypted identities are handled by the consortium, the plaintext origin of the traffic is only communicated to the requesting LEA. Therefore, the consortium members only share the final decryption shares with the LEA. Thus, no member of the consortium learns this identity. This restriction maximizes privacy, as only the requesting LEA learns the client’s identity, while the consortium only knows the relays.
The LEA then still has to verify the correctness of the decrypted client identity to ensure that the entry relay did not manipulate this information. Therefore, the LEA first verifies that the IP certificate is valid and that a trusted CA issued it. Additionally, the LEA verifies the correctness of the hop key , analogously to this verification by the consortium as previously discussed. To this end, each consortium member shares the hop key claimed in the previous layer with the LEA along its decryption share. If both checks succeed, the revealed identity is indeed that of the user from whom the suspicious traffic originated. Otherwise, the entry relay attempted to cover up the true identity, and the LEA should investigate this. Through Seldom’s iterative process, the requested identity is thus only revealed to the requesting LEA, while misbehavior (i.e., sending no or false data) during the anonymization process can be clearly attributed to one relay.
@@ 6.4. Probing Relay Honesty
Deanonymizations reveal misbehaving relays. Such misbehavior could occur due to collusion with the client, general malice toward the network, or from genuine errors. Therefore, the relay’s protocol abidance under normal circumstances should be verified without risking individuals’ privacy, such that misbehaving nodes can be fixed or removed from the network. Regular probing by the directory authorities reduces the likelihood that an exit relay’s key loss in a critical situation is coincidental. This feature significantly strengthens our assumption that failures during deanonymization are due to a relay colluding with a client, warranting further investigation.
@@
To generate probe traffic for verifying the honesty of relays, Seldom relies on cryptographic commitments. We use Pedersen commitments (Pedersen, 1991), but other commitment schemes could also be employed. To create a commitment, a directory authority, i.e., the prober, first generates a key later used as the client’s identification key during the probe. Then, the prober hashes this key, the probed exit relay, a traffic destination, and a timeframe during which the probe is active. We refer to the resulting digest as probe . To then create a Pedersen commitment to the probe , the prober generates a random value and computes the commitment on an elliptic curve with generators and .
The prober then invokes a smart contract on the translucent ledger to store . Afterward, the prober establishes a new circuit with the probed relay as the last onion router and sends traffic as specified in the commitment and using the pre-generated client identification key. Afterward, the prober opens its commitment and discloses the preimage of on the ledger. Then, the prober performs a deanonymization request, using the probe as justification for the deanonymization. The consortium accepts this request, and the exit relay is asked to initiate the deanonymization process.
@@
At this point, the exit relay does not know that the deanonymization request is only for probe traffic. It learns that the vote for the request was successful but does not see the reasoning behind it, which is only revealed during the delayed disclosure of the request to the public. The exit relay thus is expected to cooperate like for ordinary deanonymization. Afterward, the consortium verifies the signature of the database key with the public key indicated in the commitment. If this verification is successful, the exit relay passed the probe. Otherwise, the directory authority no longer accepts this exit relay. The other directory authorities can adopt this assessment but are encouraged to verify it through independent probes. Otherwise, a single directory authority would have the power to remove relays at will by claiming they failed a probe. However, a malicious relay may try to detect such additional probes and behave honestly for them to not be removed from the network. Hence, once suspicious, a relay should be probed with a higher frequency at random intervals to ensure any misbehavior was indeed a one-off event. The consortium can also further investigate the encrypted identity to identify non-exit relays that are misbehaving. Through regular probing, the consortium and the LEA can be much more confident that each relay behaves correctly during deanonymization.
In general, there are two concerns that dictate the selection of a sensible probing frequency: strain on the network and detectability. The exit bandwidth consumption of one probe is dominated by the size of the accessed web page, where a median web page is in size (http archive, 2024). Having each directory authority probe each exit relay once per day would thus result in not even of exit bandwidth consumption. To put this overhead into perspective, we can look at bandwidth authorities (AlSabah and Goldberg, 2016), which consume orders of magnitude more bandwidth for relay capacity estimations (Darir et al., 2022). Moreover, to avoid that probes are detected, the respective deanonymization needs to take time to match the manual decision-making within the consortium. Given that executing multiple probes in parallel likely adds limited value, a probing frequency above one probe per day is unrealistic.
@@
Overall, probing is expected to make up the bulk of all requests. As the public is informed about each request immediately, all probers are encouraged to reveal that a probe took place immediately after conclusion. Therefore, they can call a special function in the smart contract that triggers an event to inform the public, such that the public always knows how many requests actually deanonymize a user.
@@ 7.1. Evaluation Setup
Most new Tor features are evaluated through simulations. However, simulation abstracts from Tor’s cryptographic processing (Jansen and Hopper, 2012), which is precisely the part impacted by Seldom (and not the behavior of the network). Thus, we evaluate Seldom in a local deployment. Since our changes do not impact the network as a whole, but rather individual relays or circuits, this still gives accurate results on Seldom’s overhead.
We modify Tor version 0.4.1.5 to implement Seldom. As our target host for connection establishments, we run an apache2 web server. We additionally host a local CA for IP certificates and a MongoDB database to store deanonymization information. The consortium members’ behavior is expressed in Python, and the translucent ledger is a permissioned Ethereum blockchain (Wood and others, 2019) modified to realize the translucency properties. Our server contains two Intel® Xeon® Silver 4116 processors (24 CPU cores @ ) and of RAM. We run nine relays: one directory authority, three exit relays, and five non-exit relays. Each process (e.g., databases, Tor clients, Tor relays) is pinned to one CPU core to prevent cross-process interference.
@@ 7.2. Circuit Establishment
We compare the latency to extend a circuit by one hop at a time in Figure 6. The x-axis shows the elapsed time, while the different horizontal bars represent circuits of various lengths in Seldom or Tor, respectively. The cumulative bars represent the total time to establish a circuit.
We measure a processing overhead to extend the circuit by one hop that increases from for the first hop to for the sixth hop. This increase arises from the larger encrypted identities of longer circuits. Overall, we observe a resulting latency increase of to , depending on circuit length. The receipt of an IP certificate causes an overhead of . The remaining overhead stems from signature processing, threshold encryption, and transferring the partially encrypted identity to the next hop. All these steps require more time the more previous hops exist (as more data has to be processed). The observed variance in timings results from occasional failures of the path selection algorithm due to the small network size. In a global overlay network, additional latency in the range of multiple 100 milliseconds and jitter would be added by the network and make this overhead insignificant. Overall, Seldom thus does not perceivably impact circuit establishment.
@@ 7.3. Connection Establishment
Analyzing the impact of Seldom on connection establishment is essential, as additional latency in this step directly affects the user experience. To assess this overhead, we request a web page from the locally hosted web server through cURL, which uses socks5 to communicate with the Tor client. We measure the time to first byte, i.e., the time between issuing the request and receiving the first byte through the established circuit. We repeat this measurement for different circuit lengths (approx. 700 measurements each) and compare a vanilla Tor deployment to Seldom. Our results are shown in Figure 7. We observe that Seldom takes, on average, longer than Tor to establish a connection and receive the first byte of the response, independent of the circuit length. This delay stems from the generation and verification of the flow record signature as well as writing this flow record to a local database before the connection is established. Again, these delays are overshadowed by the latency of the overlay network, which is two orders of magnitude higher.
@@ 7.4. Bandwidth and Storage Requirements
While Seldom barely impacts performance, the additional data that is transferred and stored risks becoming the real bottleneck. To quantify this overhead, we first measure how much bandwidth exit relays consume to transfer all deanonymization data to a state-sponsored database. During our evaluation, the client establishes circuits with a length of three and a fixed exit relay. A fixed number of connections is established to random IP addresses and ports through each circuit before it is discarded. The exit relay waits for a predefined amount of closed circuits before it migrates the combined local storage for these circuits to the state-sponsored database. When migrating, mongodump is executed on the local MongoDB database, and the resulting folder is serialized and compressed using zlib. The resulting archive is sent over a TLS channel to a state-sponsored database. Simultaneously, its hash is sent to the translucent ledger to ensure consistency across multiple governmental databases.
@@
In Figure 8, we show how bandwidth usage depends on the number of connections established per circuit and how often the local buffer is flushed. Flow records are approximately an order of magnitude smaller than encrypted identities, which are only generated for each circuit and not each connection. Furthermore, the more data is migrated to the database in one go, the more the local buffer can be compressed. However, the benefit of sending more data at once levels off quickly, indicating that frequent flushing is desirable. While using an established circuit for more connections is beneficial, such a change directly influences the susceptibility to deanonymization attacks. Recent measurements by Mani et al. estimate the average number of established connections per circuit to be about 20 (Mani et al., 2018). An exit relay that migrates its database after 100 circuits have been established and used to establish 20 connections each would thus produce of data per connection.
@@
The measurements of Mani et al. also allow us to predict the storage requirements of a Tor-sized Seldom network. For this prediction, we project the amount of generated deanonymization data based on the activity in the Tor network over time. According to Mani et al., an average of 2.1 billion (95%-CI:[1.7 billion; 2.5 billion]) exit streams were created every 24 hours in April 2018 (Mani et al., 2018). Over the same period, the Tor Project reported an average of of consumed bandwidth by the Tor network (The Tor Project, 2024). For our estimations, we assume a constant relationship between the amount of consumed bandwidth and the amount of exit streams over the last six years. This is a worst-case assumption, as the size of web pages grew by over , while the number of TCP connections per web page shrunk by about over the same period (http archive, 2024). Thus, in practice, the additional bandwidth consumed by Tor is, in part, caused by larger web pages that are downloaded over fewer connections.
@@
Figure 9 thus quantifies the bandwidth overhead of Seldom in a Tor-sized network according to those estimations. Here, the blue line shows the global bandwidth consumed by all exit relays to write data into the state-sponsored database over time with confidence intervals (blue area). In January 2020, all exit relays combined would consume (95%-CI:[; ]) to write deanonymization data to the state-sponsored database. To put these numbers into perspective, this would mean a mere (95%-CI:[; ]) increase in data transmitted by exit relays compared to Tor. A Seldom exit relay advertising of exit traffic per second would thus insert approx. of deanonymization data into the database. This estimate also means that this same relay has to, on average, verify 385 ed25519 signatures and compute 18 hybrid threshold encryptions per second, a processing overhead hardly impacting exit relays. Figure 9 also highlights the amount of data that would be retained in the state-sponsored database, assuming a retention period of two years. The database would thus have to store (95%-CI:[; ]) of deanonymization data, which can easily be provided by government organizations.
@@
Seldomcan thus easily scale to Tor’s current size. If Seldom were to significantly outgrow Tor, there exist several strategies to improve scalability. First, websites that are unlikely to be abused for illegal activities (e.g., news websites) could be whitelisted, i.e., exit relays would not store flow records (and encrypted identities if only such websites are visited). Secondly, Seldom could rely on probabilistic deanonymization, where only a randomly selected subset of identities is deanonymizable. Thus, the consortium could retroactively select a fraction of records that are persistently stored through provable randomness. Thus, bandwidth and storage overhead are unlikely to limit Seldom’s scalability.